trish:

Are the following lines parallel, perpendicular, or neither? 3x + 5y = 40 11 - y = 3/5x

Hero:

@trish, ask yourself the following question: Are the slopes of both lines the same or are they negative reciprocals of one another?

trish:

@hero yes they are

trish:

@hero oh so it's perpendicular

Hero:

@trish, Why do you think they are perpendicular?

Hero:

And can you please verify your email? You will not be able to use the messaging system unless you do so.

trish:

@hero it's not sending me an email

trish:

@hero I thought they were perpendicular because of the slope being flipped first its 3,5. Then in the next line it has 3/5

Hero:

You're not receiving any emails because you have not verified the initial email that was sent to you.

Hero:

You have to check both your spam and your inbox to find it.

trish:

@hero ok I verified it so about the question am I right

Hero:

Okay so here are the two lines: 3x + 5y = 40 11 - y = (3/5)x Be sure to use parentheses so that people do not interpret \(\dfrac{3}{5}x\) to mean \(\dfrac{3}{5x}\).

trish:

@Hero so it is perpendicular?

Hero:

So the thing is if you multiply both sides of the 2nd equation by 5 you end up with 55 - 5y = 3x Then adding 5y to both sides you get 3x + 5y = 55 which shows that the lines both have the same slope.

trish:

are they parallel?

Hero:

What are your thoughts regarding this? Do you understand what I did?

trish:

@hero no

Hero:

So the 2nd equation gave you was \(11 - y = \dfrac{3}{5}x\) correct?

Hero:

@trish

trish:

yes @hero

Hero:

Okay so what happens if you try multiplying both sides by 5?

trish:

that cancels out and it leaves you with 55-5y=3x

trish:

@Hero

Hero:

Yes correct. And then as I said, when you add 5y to both sides, you end up with 3x + 5y = 55. Compare that with 3x + 5y = 40.

trish:

@hero so that means it's neither

Hero:

Have you tried graphing them on http://www.desmos.com?

trish:

@hero I'll try

Hero:

:D

trish:

@Hero oh ok they're parallel

trish:

@hero thanks for your help

Hero:

Yes, parallel correct.

Hero:

@trish